Each letter in the following multiplication problem represents a unique digit. Determine the letters that would make the multuplication problem correct:
1PANSY
x 3
_______
PANSY1

Answer by Edwin McCravy(20067)

Let x = PANSY

1PANSY = 100000 + PANSY = 100000 + x

PANSY1 = 10*PANSY + 1 = 10x + 1

Since 1PANSY*3 = PANSY1,

    (100000 + x)*3 = 10x + 1

     3(100000 + x) = 10x + 1

       300000 + 3x = 10x + 1

            299999 = 7x

 Divide both sides by 7

             42857 = x

So PANSY = 42857

So

1PANSY = 142857

PANSY1 = 428571   

Checking

142857*3 = 428571

Edwin